OpenStudy (unklerhaukus):

consider the differential equation \[y''+\lambda y = 0\] \[y(0) = 0 , y(π) = 0\] find the eigenvalues and corresponding eigenfunctions, @mathematics

OpenStudy (jamesj):

What are you finding hard about this? Apart from the eigensomething language, you've solved this problem a few dozen times by now, haven't you?

OpenStudy (unklerhaukus):

well i have forgotten how to begin, and how to finish, Ill probably be able to do the mathematics but i am struggling to turn the question into equations

OpenStudy (jamesj):

Consider three cases in turn. lambda > 0, = 0, < 0. Find the general solution for each such lambda; see if it can be made to meet the initial conditions. The lambda for which you can satisfy those initial conditions are the eigenvalues and the corresponding solutions the eigenfunctions.

OpenStudy (jamesj):

btw, you're kind of stingy in acknowledging help. Not that I'm neurotically checking, really I'm not, but the occasional medal for my amazing insights would be appreciated. ;-)

OpenStudy (unklerhaukus):

What are you talking about i always reward a good answer with a medal.

OpenStudy (jamesj):

Ok then. Anyway, analyze this as I suggest. Basically what you're doing is finding the eigenvectors of the operator \( D^2 \) acting on the \( C^\infty(\mathbb{R}) \). Lambda here is the negative of the eigenvalues. I.e., if nu is an eigenvalue, then \[ D^2 = \nu \iota \] where iota is the identity function, and thus \( \lambda = -\nu \).

OpenStudy (jamesj):

The set of eigenvalues an operator such as D^2 has is called the spectrum of the operator. And once you learn to think about this property of operators, it opens up rooms and rooms of insight.

OpenStudy (unklerhaukus):

well i am currently fresh outta insight and stocked up on confusion and malaise i will do this problem someothertime